Suppose a hypothetical reaction A ---------> D + E

has a t1/2 of 8.50 min when the initial concentration of A is 0.150 M.

How long will it take for the concentration to drop to 0.035 M if the reaction is first order with respect to A.

Please write down the steps

Thank you
Answer #1

Half life time t​​​​​​1/2 = 8.50 min .......(given)

Initial concentration (C​​​​​​0) = 0.150 M .....(given)

Reduced concentration (C​​​​​​t) = 0.035 M ....(given)

Time =?

We know, half life time for a first order reaction is calculated by formula -

t​​​​​​1/2 = 0.693/K

Where, K = rate constant ; t​​​​​​1/2 half life time

Or K = 0.693/t​​​​​​1/2

So, k = 0.693/8.50 min

= 0.0815 min^-1

First order reaction is calculated by formula -

t = 1/K × ln (C​​​​​​0 / C​​​​​​t)

t = 1/ 0.0815 × ln (0.150 M/0.035 M)

= 12.269 min × 1.455

= 17.85 min.

Hence time it will take for the concentration reduced to 0.035 M is = 17.85 min.

Please rate the question. Thank you!